OpenStudy (anonymous):

Given x/12=5/4 , which is proportional to x+5/x-5

OpenStudy (anonymous):

Given x/12 = 5/4, which is proportional to x+5/x-5 ( The / means it's a fraction ) .x . . 5 --- = ---- 12 . .4 4x + 5(12) 4x = 60 x = 15 x + 5 -------- = x - 5 15 + 5 ---------- = 15 - 5 20 . . .2 ---- = ---- 10 . . .1 - 2/1 <==ANSWER